OpenStudy (anonymous):

-2r(8r+5) A. -16r^2 + 5 B. -16r^2 +10 C. -16r^2 -10r D. -16r^2 -10

jigglypuff314 (jigglypuff314):

distributive property works like a(b+c) = ab + ac so -2r(8r+5) = (-2r)(8r) + (-2r)(5) multiply and simplify :)

OpenStudy (anonymous):

nicely explained :)

OpenStudy (anonymous):

I don't get it? @jigglypuff314

jigglypuff314 (jigglypuff314):

hmmm what do you not understand? :)

OpenStudy (anonymous):

how to solve the problem

jigglypuff314 (jigglypuff314):

well I explained how to do the first step, do you understand that? :)

OpenStudy (anonymous):

no :/

jigglypuff314 (jigglypuff314):

It's like the idea that you want to give the outside number to the inside numbers for example, 3(x+8) you would give the 3 to the x and then you would give the 3 to the 8 so = 3x + (3)(8) = 3x + 24 do you understand that? :)

OpenStudy (anonymous):

yes

jigglypuff314 (jigglypuff314):

then it's the same idea for -2r(8r+5) you want to give (-2r) to the 8r and the (-2r) to the 5 so you get (-2r)(8r) + (-2r)(5)

OpenStudy (anonymous):

yes thanks

jigglypuff314 (jigglypuff314):

then the step after that would be to multiply like (-2r) times (8r) = (-2times8)r² = -16r²

OpenStudy (anonymous):

I got A. after I solved the problem is this right? :D

jigglypuff314 (jigglypuff314):

hmmm not quite :) (-2r)(5) = -2r times 5 = (-2 times 5) r = ?

OpenStudy (anonymous):

-10

jigglypuff314 (jigglypuff314):

don't forget the r :) so -10r